At least one of the answers above is NOT correct. 2 1 (1 pt) Let M 2 5 Find formulas for the entries of M, where n is a posican someone help me with this~
Answer #1

2 Sa Gun madne M Claractennhe quthan in M-AT]=o 2-7 - (2--2 (5)+20 1o-sa-2rtt+ 2=0 Xニ1+12 =0 t47-32412=0 Hn diagaalizable (A-